OpenStudy (anonymous):

A balloon has a circumference of 22 in. Use the circumference to approximate the surface area of the balloon to the nearest square centimeter.

OpenStudy (oaktree):

Balloons don't have a circumference.

OpenStudy (campbell_st):

you need to use the circumference to find the radius... \[C=2\pi r\] \[22 = 2\pi r\] \[\frac{22}{2\pi} = r\] \[r \approx 3.5\] next the surface area of a sphere is \[SA = 4\pi r^2\] JUts substitute r and evaluate